OpenStudy (anonymous):

find exact value of arctan(-1)

OpenStudy (anonymous):

y=arctan(-1) tany=-1 y=3pi/4

OpenStudy (anonymous):

same type of question

OpenStudy (anonymous):

but use the CAST rule

OpenStudy (anonymous):

okies

OpenStudy (anonymous):

tyvm :)

OpenStudy (anonymous):

np, gonna sleep

OpenStudy (anonymous):

gn sleep well

OpenStudy (nikvist):

\[\arctan{x}:\quad\mathbb{R}\rightarrow[-\pi/2,\pi/2]\]\[\Rightarrow\quad\arctan{(-1)}=-\pi/4\]